# Learning classical density functionals for ionic fluids

### v1

@@ Abstract
Accurate and efficient theoretical techniques for describing ionic fluids are highly desirable for many applications across the physical, biological and materials sciences. With a rigorous statistical mechanical foundation, classical density functional theory (cDFT) is an appealing approach, but the competition between strong Coulombic interactions and steric repulsion limits the accuracy of current approximate functionals. Here, we extend a recently presented machine learning (ML) approach [Sammüller et al., Proc. Natl. Acad. Sci. USA, 120, e2312484120 (2023)] designed for systems with short-ranged interactions to ionic fluids. By adopting ideas from local molecular field theory, the framework we present amounts to using neural networks to learn the local relationship between the one-body direct correlation functions and inhomogeneous density profiles for a “mimic” short-ranged system, with effects of long-ranged interactions accounted for in a mean-field, yet well-controlled, manner. By comparing to results from molecular simulations, we show that our approach accurately describes the structure and thermodynamics of the prototypical model for electrolyte solutions and ionic liquids: the restricted primitive model. The framework we present acts as an important step toward extending ML approaches for cDFT to systems with accurate interatomic potentials.
The behavior of ionic fluids underlies a vast array of physical and biological phenomena as well as technological applications, ranging from electrolyte solutions controlling protein folding [1] to room-temperature ionic liquids in energy storage devices [2]. A fundamental topic that continues to attract enormous interest both experimentally [3, 4, 5, 6, 7] and theoretically [8, 9, 10, 11, 12, 13] is the structure and thermodynamics of ions near charged interfaces, in particular, how the nature of both the electrolyte and solid surface impacts the properties of the electric double layer (EDL). Poisson–Boltzmann (PB) theory and its linearized Debye-Hückel form provide the basis for much of our understanding of ionic fluids. Their neglect of correlations arising from nonelectrostatic interactions, however, restricts their validity to fluids of low ionic strength.
@@
Two key features underpin the success of the neural functional approach: (i) the feasibility of GC simulations to produce inhomogeneous density profiles of sufficient quality to form a reliable training set; and (ii) short-ranged (SR) correlations such that the functional relationship between $c_{\nu}^{(1)}$ and $\{\rho_{\nu}\}$ is local. Ionic fluids pose challenges on both fronts. First, GC simulations for ionic systems raise delicate questions concerning electroneutrality, with various schemes proposed that either insert individual ions or neutral pairs [38, 39, 40, 41, 42]. Even then, simulations corresponding exactly to the GC ensemble where the number of each species can fluctuate are impractical, owing to poor acceptance rates of trial insertion and deletion moves. Second, the long-ranged (LR) nature of the Coulomb potential leads to a nonlocal relationship between $c_{\nu}^{(1)}$ and $\{\rho_{\nu}\}$. We circumvent both of these issues by adopting concepts from local molecular field theory (LMFT) [43, 44, 45, 46, 47], which has been shown to have close links to cDFT [48].
We focus our efforts on the protypical model of an ionic fluid: the restricted primitive model (RPM) comprising oppositely charged hard spheres of equal diameter $\sigma$, embedded in a uniform dielectric continuum. The scheme we propose can be briefly summarized. First, by employing neural networks, we find the one-body direct correlation functions for a suitably chosen “mimic system” whose electrostatic interactions are entirely short-ranged. Then, by leveraging LMFT and its relation to cDFT, we account for the net averaged effects of LR electrostatic interactions in a well-controlled fashion. When compared to molecular simulations, the framework we outline describes inhomogeneous density profiles, the equation of state, and the properties of the electric double layer in the presense of electric fields with very high accuracy.
@@
To learn the functional relationship for $c^{(1)}_{\mathrm{R},\nu}(\bm{\mathit{r}};[\{\rho_{\mathrm{R},\nu}\}])$, we obtain data for the right hand side of Eq. 3 by measuring $\rho_{\mathrm{R},\nu}$ from simulations with known $\beta V_{\mathrm{R},\nu}$, $\beta\phi_{\mathrm{R}}$ and $\beta\mu_{\mathrm{R},\nu}$. To this end, in line with Ref. 36, we perform simulations in a planar geometry at different combinations of $\{\beta V_{\mathrm{R},\nu}\}$, $\beta\phi_{R}$ and $\{\beta\mu_{\mathrm{R},\nu}\}$. The form of $v_{0}$ means that each particle of the mimic system can be considered electroneutral, comprising both a point charge and a compensating Gaussian charge distribution [51]. As such, in addition to translational moves, we can readily perform GC particle insertions/deletions, along with semi-GC swapping of “anions” and “cations” without needing to worry about issues of electroneutrality (see Fig. 1(a)). We have found this highly beneficial to converging our simulations, full details of which are provided in the SM [49].
For each species, $\nu=+\,\mathrm{or}\,-$, the neural network used to represent the relationship $\{\rho_{\mathrm{R},\nu}(z)\}\rightarrow c^{(1)}_{\mathrm{R},\nu}(z)$ is structured as follows. The input layer has two channels which are supplied with the discretized values of the cation and anion density profiles in a window $\Delta z=3.6\,\sigma$ around the particular value of $z$. Following a fully-connected multilayer perceptron of three layers, the output layer consists of a single node, which yields the predicted value of $c^{(1)}_{\mathrm{R},\nu}$ at position $z$. For the SR mimic RPM, we train two independent networks, one for the cation and one for the anion. To reduce noise in the bulk structure predictions, we also obtained models regularized with bulk two-body direct correlation functions as proposed in Ref. 52. Details of our training procedure are provided in the SM [49], and all data are openly available.
Figure 2: Structure and thermodynamics of the LR full system. cDFT for the RPM at $T^{*}=0.066$ with effects of LR electrostatic interactions accounted by LMFT shows excellent agreement with reference molecular simulation data for (a) the equation of state, (b) ion density profiles confined in a slit and (c) with an applied external potential. In (a), we see that our cDFT approach outperformcs the prediction obtained by combining the Carnahan–Starling equation of state for hard-sphere fluids ($P^{\mathrm{CS}}$) and the mean spherical approximation ($P^{\mathrm{MSA}}$). For the LR system in (c), the applied external potential acts over a much longer range and the chemical potentials are shifted lower compared to the SR mimic system in Fig. 1(c) to yield $\rho_{\nu}(z)=\rho_{\mathrm{R},\nu}(z)$.
@@
Fig. 1(b) shows that the static structure factors obtained from the functional are in excellent agreement with results from a molecular dynamics (MD) simulation of the mimic system at the same bulk densities. The symmetry of the RPM is well-captured, reflected in $S_{NZ}(k)=0$. Moreover, by comparing to results from a MD simulation of the full system, we see that $S_{NN}(k)$ and $S_{ZZ}(k)$ for the full and mimic systems agree very well at sufficiently large $k$, confirming that our choice of $\kappa^{-1}=1.8\,\sigma$ is sufficient for the mimic system to capture the SR correlations of the full system. Deviations of the SR system from the LR system only manifest significantly in $S_{ZZ}(k)$ at small $k$, in agreement with previous works on the subject [46, 55]. In particular, as shown in the inset, we see that $S_{ZZ}(k)$ for the LR system strictly obeys the Stillinger–Lovett sum rule [56, 57]
@@
$\begin{split}\Lambda^{3}_{\nu}\rho_{\mathrm{R},\nu}(\bm{\mathit{r}})=\exp\Big{(}-\beta V_{\mathrm{R},\nu}(\bm{\mathit{r}})&-\beta q_{\nu}\phi_{R}(\bm{\mathit{r}})\\ &+\beta\mu_{\mathrm{R},\nu}+c^{(1)}_{\mathrm{R},\nu}(\bm{\mathit{r}};[\{\rho_{\nu}\}])\Big{)},\end{split}$ (6)
@@
$c_{\nu}^{(1)}(\bm{\mathit{r}};[\{\rho_{\nu}\}])=c_{\mathrm{R},\nu}^{(1)}(\bm{\mathit{r}};[\{\rho_{\nu}\}])-\beta\Delta\mu_{\nu}-\beta q_{\nu}\Delta\phi(\bm{\mathit{r}}),$ (8)
@@
$\Delta\phi(\bm{\mathit{r}})\equiv\phi(\bm{\mathit{r}})-\phi_{\mathrm{R}}(\bm{\mathit{r}})=-\frac{1}{\epsilon}\int\!\!\mathrm{d}\bm{\mathit{r}}\,n(\bm{\mathit{r}})\,v_{1}(|\bm{\mathit{r}}-\bm{\mathit{r}}^{\prime}|),$ (9)
@@
Based on Ref. 59, we have further derived a correction for the chemical potential (see the SM [49]),
@@
The electric double layer. The results presented so far demonstrate that our cDFT approach provides an efficient and accurate route to the structure and thermodynamics of bulk and inhomogenous ionic fluids. As an application, we turn our attention to the fundamental topic of ongoing scientific interest: the electric double layer.
As a simple model of an EDL-forming system, we consider the RPM confined between two repulsive walls, and in the presence of an electric field $E_{z}$ along the $z$ direction. The electrostatic potential that the LR system feels is $\phi(z)=-E_{z}z$, as depicted in Fig. 3(a). As pointed out in previous applications of LMFT [63], the SR mimic system in this case feels the electrostatic potential $\phi_{\mathrm{R}}$ due to smooth planar Gaussian charge distributions at the walls, as depicted in Fig. 3(b). For the reference simulation data, the LR system was simulated at constant $E_{z}$ using the finite-field approach [64, 65, 66, 67, 68, 69, 70]. As shown in Fig. 3(c), the ion distributions from our cDFT approach are in excellent agreement with the results from simulation. As a significant advancement in theoretical descriptions of the EDL, our cDFT approach can accurately describe the strong oscillations in the density profiles of co- and counter-ions. Moreover, as we detail in the SM [49], our functional is thermodynamically consistent, satisfying the contact density theorem [71, 72, 73]
@@
which relates the bulk pressure to the surface charge density $\sigma_{\rm s}=\int_{-L}^{0}\!\mathrm{d}z\,n(z)$  [^3] This equation for $\sigma_{\rm s}$ assumes that the electrolyte is centered in a periodically replicated cell of length $L$ along the $z$ direction. and ion densities at contact with the walls. Obeying this contact theorem has proven a challenge for most state-of-the-art cDFT functionals for ionic fluids [75, 76].
Figure 3: Accurately describing the EDL with cDFT. For the RPM confined between two repulsive walls and in the presence of an electric field, the potentials experienced by the full and mimic systems are shown in (a) and (b), respectively. The resulting ion density profiles from cDFT are in excellent agreement with molecular simulations, and obey the contact theorem (Eq. 15).
The main advantage of the neural functional approach outlined in Ref. 36 is arguably that the local relationship between $c_{\nu}^{(1)}$ and $\{\rho_{\nu}\}$ permits application of the resulting functional to system sizes far beyond those encountered during training of the ML. By using the framework of LMFT and its relationship to cDFT, we have successfully extended the neural functional approach to a case where the relationship between $c_{\nu}^{(1)}$ and $\{\rho_{\nu}\}$ is nonlocal. Future work will pursue using neural functionals to understand EDL capacitance [77] and surface force balance measurements [13], and to augment other cDFT approaches to understand solvation [78]. The framework we have outlined should also find use in describing the dielectric response of polar fluids such as water [79, 80, 55].

### v2

@@ Abstract
Accurate and efficient theoretical techniques for describing ionic fluids are highly desirable for many applications across the physical, biological and materials sciences. With a rigorous statistical mechanical foundation, classical density functional theory (cDFT) is an appealing approach, but the competition between strong Coulombic interactions and steric repulsion limits the accuracy of current approximate functionals. Here, we extend a recently presented machine learning (ML) approach [Sammüller et al., Proc. Natl. Acad. Sci. USA, 120, e2312484120 (2023)] designed for systems with short-ranged interactions to ionic fluids. By adopting ideas from local molecular field theory, the framework we present amounts to using neural networks to learn the local relationship between the one-body direct correlation functions and inhomogeneous density profiles for a “mimic” short-ranged system, with effects of long-ranged interactions accounted for in a mean-field, yet well-controlled, manner. By comparing to results from molecular simulations, we show that our approach accurately describes the structure and thermodynamics of prototypical models for electrolyte solutions and ionic liquids, including size-asymmetric and multivalent systems. The framework we present acts as an important step toward extending ML approaches for cDFT to systems with accurate interatomic potentials.
The behavior of ionic fluids underlies a vast array of physical and biological phenomena as well as technological applications, ranging from electrolyte solutions controlling protein folding [1] to room-temperature ionic liquids in energy storage devices [2]. A fundamental topic that continues to attract enormous interest both experimentally [3, 4, 5, 6, 7] and theoretically [8, 9, 10, 11, 12, 13] is the structure and thermodynamics of ions near charged interfaces, in particular, how the nature of both the electrolyte and solid surface impacts the properties of the electric double layer (EDL). Poisson–Boltzmann (PB) theory and its linearized Debye-Hückel form provide the basis for much of our understanding of ionic fluids. Their neglect of correlations arising from non-electrostatic interactions, however, restricts their validity to fluids of low ionic strength.
@@
Two key features underpin the success of the neural functional approach: (i) correlations in the fluid are short-ranged (SR), such that the functional relationship between $c_{\nu}^{(1)}$ and $\{\rho_{\nu}\}$ is local; and (ii) the feasibility of GC simulations to produce inhomogeneous density profiles of sufficient quality to form a reliable training set. Ionic fluids pose challenges on both fronts. First, the long-ranged (LR) nature of the Coulomb potential leads to a nonlocal relationship between $c_{\nu}^{(1)}$ and $\{\rho_{\nu}\}$. Second, GC simulations for ionic systems raise delicate questions concerning electroneutrality, with various schemes proposed that either insert individual ions or neutral pairs [38, 39, 40, 41, 42]. Even then, simulations corresponding exactly to the GC ensemble where the number of each species can fluctuate are impractical, owing to poor acceptance rates of trial insertion and deletion moves. We circumvent both of these issues by adopting concepts from local molecular field theory (LMFT) [43, 44, 45, 46, 47], which has been shown to have close links to cDFT [48].
We initially focus our efforts on the prototypical model of an ionic fluid: the restricted primitive model (RPM) comprising oppositely charged hard spheres of equal diameter $\sigma$, embedded in a uniform dielectric continuum. Later, we will also present results for a primitive model (i.e., size-asymmetric) and a multivalent system. The scheme we propose can be briefly summarized. First, by employing neural networks, we find the one-body direct correlation functions for a suitably chosen “mimic system” whose electrostatic interactions are entirely short-ranged. Then, by leveraging LMFT and its relation to cDFT, we account for the net averaged effects of LR electrostatic interactions in a well-controlled fashion. When compared to molecular simulations, the framework we outline describes inhomogeneous density profiles, the equation of state, and the properties of the electric double layer in the presense of electric fields with very high accuracy.
@@
To learn the functional relationship for $c^{(1)}_{\mathrm{R},\nu}(\bm{\mathit{r}};[\{\rho_{\mathrm{R},\nu}\}])$, we obtain data for the right hand side of Eq. 3 by measuring $\rho_{\mathrm{R},\nu}$ from simulations with known $\beta V_{\mathrm{R},\nu}$, $\beta\phi_{\mathrm{R}}$ and $\beta\mu_{\mathrm{R},\nu}$. To this end, in line with Ref. 36, we perform GC simulations in a planar geometry at different combinations of $\{\beta V_{\mathrm{R},\nu}\}$, $\beta\phi_{R}$ and $\{\beta\mu_{\mathrm{R},\nu}\}$. Note that GC simulations are essential for the purpose of evaluating Eq. 3 and, on their own, canonical methods such as molecular dynamics (MD) simulations are insufficient. The form of $v_{0}$ means that each particle of the mimic system can be considered electroneutral, comprising both a point charge and a compensating Gaussian charge distribution [51]. As such, in addition to translational moves, we can readily perform GC particle insertions/deletions, along with semi-GC swapping of “anions” and “cations” without needing to worry about issues of electroneutrality [see Fig. 1(a)]. We have found this highly beneficial to converging our simulations, full details of which are provided in the SM [49]. In total, $\sim 2500$ simulations have been performed to gather training data.
For each species, $\nu=+\,\mathrm{or}\,-$, the neural network used to represent the relationship $\{\rho_{\mathrm{R},\nu}(z)\}\rightarrow c^{(1)}_{\mathrm{R},\nu}(z)$ is structured as follows. The input layer has two channels that are supplied with the discretized values of the cation and anion density profiles in a window $\Delta z=3.6\,\sigma$ around the particular value of $z$. Following a fully-connected multilayer perceptron of three layers, the output layer consists of a single node, which yields the predicted value of $c^{(1)}_{\mathrm{R},\nu}$ at position $z$. For each ionic fluid considered, we train two independent networks, one for the cation and one for the anion. To reduce noise in the bulk structure predictions, we also obtained models regularized with bulk two-body direct correlation functions as proposed in Ref. 52. Details of our training procedure are provided in the SM [49], and all data are openly available [53].
Figure 2: Structure and thermodynamics of the LR full system. cDFT for the RPM at $T^{*}=0.066$ with effects of LR electrostatic interactions accounted by LMFT shows excellent agreement with reference molecular simulation data for (a) the equation of state, (b) ion density profiles confined in a slit and (c) with an applied external potential. In (a), we see that our cDFT approach outperforms the prediction obtained by combining the Carnahan–Starling equation of state for hard-sphere fluids ($P^{\mathrm{CS}}$) and the mean spherical approximation ($P^{\mathrm{MSA}}$). For the LR system in (c), the applied external potential acts over a much longer range and the chemical potentials are shifted lower compared to the SR mimic system in Fig. 1(c) to yield $\rho_{\nu}(z)=\rho_{\mathrm{R},\nu}(z)$.
@@
Fig. 1(b) shows that the static structure factors obtained from the functional are in excellent agreement with results from a molecular dynamics simulation of the mimic system at the same bulk densities  [^2] While GC simulations (i.e., known $\{\mu_{\nu}\}$) are required to train the neural networks (see Eq. 3), it is reasonable to compare to average structures obtained with canonical simulations at the same average density.. The symmetry of the RPM is well-captured, reflected in $S_{NZ}(k)=0$. Moreover, by comparing to results from a MD simulation of the full system, we see that $S_{NN}(k)$ and $S_{ZZ}(k)$ for the full and mimic systems agree very well at sufficiently large $k$, confirming that our choice of $\kappa^{-1}=1.8\,\sigma$ is sufficient for the mimic system to capture the SR correlations of the full system. Deviations of the SR system from the LR system only manifest significantly in $S_{ZZ}(k)$ at small $k$, in agreement with previous works on the subject [46, 57]. In particular, as shown in the inset, we see that $S_{ZZ}(k)$ for the LR system strictly obeys the Stillinger–Lovett sum rule [58, 59]
@@
$\begin{split}\Lambda^{3}_{\nu}\rho_{\mathrm{R},\nu}(\bm{\mathit{r}})=\exp\Big{(}-\beta V_{\mathrm{R},\nu}(\bm{\mathit{r}})&-\beta q_{\nu}\phi_{R}(\bm{\mathit{r}})\\ &+\beta\mu_{\mathrm{R},\nu}+c^{(1)}_{\mathrm{R},\nu}(\bm{\mathit{r}};[\{\rho_{\mathrm{R},\nu}\}])\Big{)},\end{split}$ (6)
@@
$c_{\nu}^{(1)}(\bm{\mathit{r}};[\{\rho_{\nu}\}])=c_{\mathrm{R},\nu}^{(1)}(\bm{\mathit{r}};[\{\rho_{\nu}\}])-\beta\Delta\mu_{\nu}+\beta q_{\nu}\Delta\phi(\bm{\mathit{r}}),$ (8)
@@
$\Delta\phi(\bm{\mathit{r}})\equiv\phi(\bm{\mathit{r}})-\phi_{\mathrm{R}}(\bm{\mathit{r}})=-\frac{1}{\epsilon}\int\!\!\mathrm{d}\bm{\mathit{r}}^{\prime}\,n(\bm{\mathit{r}}^{\prime})\,v_{1}(|\bm{\mathit{r}}-\bm{\mathit{r}}^{\prime}|),$ (9)
@@
Based on Ref. 61, we have further derived a correction for the chemical potential (see the SM [49]),
@@
The electric double layer and other ionic fluids. The results presented so far demonstrate that our cDFT approach provides an efficient and accurate route to the structure and thermodynamics of the RPM. As an application, we turn our attention to the fundamental topic of ongoing scientific interest: the electric double layer. We also show that our approach is robust to choice of interatomic potential. Specifically, we also consider the primitive model (PM) and a multivalent fluid.
As a simple model of an EDL-forming system, we consider the RPM confined between two repulsive walls, and in the presence of an electric field $E_{z}$ along the $z$ direction. The electrostatic potential that the LR system feels is $\phi(z)=-E_{z}z$. For the reference simulation data, the LR system was simulated at constant $E_{z}$ using the finite-field approach [65, 66, 67, 68, 69, 70, 71]. As shown in Fig. 3(a), the ion distributions from our cDFT approach are in excellent agreement with the results from simulation. Notably, the cDFT accurately describes the strong oscillations in the density profiles of co- and counter-ions. Moreover, as we detail in the SM [49], our functional is thermodynamically consistent, satisfying the contact density theorem [72, 73, 74]
@@
which relates the bulk pressure to the surface charge density $\sigma_{\rm s}=\int_{-L}^{0}\!\mathrm{d}z\,n(z)$  [^4] This equation for $\sigma_{\rm s}$ assumes that the electrolyte is centered in a periodically replicated cell of length $L$ along the $z$ direction. and ion densities at contact with the walls. Obeying this contact theorem has proven a challenge not only for integral equation methods [76, 77, 78, 79], which tend to violate key sum rules and are therefore generally thermodynamically inconsistent [80, 48], but also for most state-of-the-art cDFT functionals for ionic fluids [81, 82]. The accurate description of EDL structure and thermodynamic consistency displayed by our LMFT-based neural cDFT approach constitutes a significant advancement in the theoretical description of ionic fluids.
Figure 3: Accurately describing the EDL with cDFT for (a) the RPM, (b) the PM, and (c) a multivalent fluid. In all cases, confining walls located at ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z/\sigma=\pm 7.2}$ in a periodic cell of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L/\sigma=18.1}$, and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}E_{z}=1.7\,k_{\mathrm{B}}T/e\sigma}$. The resulting ion density profiles from cDFT are in excellent agreement with molecular simulations, and obey the contact theorem (Eq. 15).
The main advantage of the neural functional approach outlined in Ref. 36 is arguably that the local relationship between $c_{\nu}^{(1)}$ and $\{\rho_{\nu}\}$ permits application of the resulting functional to system sizes far beyond those encountered during training of the ML. By using the framework of LMFT and its relationship to cDFT, we have successfully extended the neural functional approach to a case where the relationship between $c_{\nu}^{(1)}$ and $\{\rho_{\nu}\}$ is nonlocal. To demonstrate that this methodology is not limited to the RPM, in Figs. 3(b) and (c), we present results for a PM and multivalent system. For the PM, we have changed the sizes of the anion and cation to $\sigma_{-}=4\sigma/3$ and $\sigma_{+}=2\sigma/3$, respectively. For the multivalent system, we consider equal-sized anions and cations of diameter $\sigma$, but increase the valency of the anion such that $q_{-}/q_{+}=2$. We use the same temperature as for the RPM, $T^{*}=\epsilon\sigma k_{\mathrm{B}}T/|q_{+}q_{-}|=0.066$. In both cases, we observe excellent agreement between the neural cDFT and simulations, both for inhomogeneous profiles and bulk equations of state, as further shown in the SM [49]). Future work will pursue using neural functionals to understand EDL capacitance [83] and surface force balance measurements [13], and to augment other cDFT approaches to understand solvation [84]. The framework we have outlined should also find use in describing the dielectric response of polar fluids such as water [85, 86, 57].